PLEASE HELP I AM A LITTLE BIT CONFUSED THANK YOU
The sum of two consecutive integers is 145. Find the two integers.

Answer by scott8148(6628)

consecutive integers means that they are one number apart (5,6 ; 22,23 ; etc.) so if one is x, the other is x+1

x+(x+1)=145 ... 2x+1=145 ... 2x=144 ... x=72 ... x+1=73 ... 72+73=145